Shift report 20000721 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 21 2000 - 16:02:18 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000721 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Jul 21 16:02:18 2000, Shift: 08:00-16:00
    Supervisor: PC, with: EJK
    Log book#3 pages 39 - 41
    
    Shift Summary:
    ------------------------------------------------------------
    There has been some beam at injection energy. We have a 
    run 1880 written to spool that can be used for a little debugging,
    and a large cosmic ray run (~300K) was finished this morning 
    (run 1869 ).
    Once they tried to ramp the beam but immediatly lost the 
    yellow.
    
    Noted that BRAHVO does not have HV info for PMT Top 44-83 and 
    tried to add load a correct file, but it did not change anything.
    Maybe there is some max in BRAHVO ?
    
    We have added TPCOnlineMonitor and already the first
    complaint is there. We have seen that it crash when we use
    TPC pedestal data. Some problem in BrTPCClusterFinder and 
    this is likely the problem in the other case as well.
    
    Looked at hodoscope H1 and H2. Problem was often that the 
    TDC cables for 2 nearby slats had been interchanged. One 
    bad channel in H2 ADC was due to problems in the delay cables.
    The HV problems were not fixed - needs access.
    
    Have a nice weekend
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 21 2000 - 16:05:48 EDT